Cookie bakes 25 pies for the farmers market. She charges $7 for an apple pie and $10 for a cherry pie, she sells out of all her pies so her total sales are $205.
How many pies were apple and how many were cherry?

Found 2 solutions by stanbon, cooper30:
Answer by stanbon(75887)

Cookie bakes 25 pies for the farmers market. She charges $7 for an apple pie and $10 for a cherry pie, she sells out of all her pies so her total sales are $205.
How many pies were apple and how many were cherry?
------
Quantity Equation::: a + c = 25
Value Equation:::: 7a+10c = 205
------
Modify for elimination:
7a + 7c = 7*25
7a +10c = 205
---------------------
Solve for "c":
3c = 30
c = 10 (# of cherry pies sold)
---------------------------
Solve for "":
a + c = 25
a + 10 = 25
a = $15 (# of apple pies sold)
================================
Cheers,
Stan H.
==========

Answer by cooper30(1)

Apple = 5EA
Cherry = 17